# The Impact of Natal Kicks on Black Hole Binaries

### v1

@@ 3 Characteristic Orbit of the Pre-Supernova Binary
Figure 4 presents $v_{\rm lim}$ for binaries with primary star mass between $1\leq M_{1}/M_{\odot}\leq 1000$ and mass ratios $q:=M_{2}/M_{1}=\{0.1,1\}$. The black line indicates the $v_{\rm lim}$ threshold that determines which binary BHs (formed without natal kicks) can merge within the age of the Universe; values below this line correspond to binaries that are too wide to merge and therefore cannot become gravitational-wave sources. For a given mass ratio, the region between the blue and red lines corresponds to natal kick magnitudes that, for that specific orbital configuration, always result in retrograde orbits. Finally, the red line marks the limit above which binaries become gravitationally unbound.
@@
For a star mass of $10\leavevmode\nobreak\ M_{\odot}$ with a mass ratio $q=0.1$ (or $q=1$), the limit orbital speed is $v_{\rm lim}\approx 41$ km/s (or $v_{\rm lim}\approx 213$ km/s). This implies that to disrupt the binary, the natal kicks must exceed roughly $100$ km/s (or $515$ km/s). It is important to note that progenitors of binary BH mergers may exhibit higher orbital speeds, which would in turn raise these limit values. Therefore, our current estimates for the limiting orbital speeds can be considered conservative. In the context of the complete collapse scenario, most models suggest that neutrino natal kicks are $\sim 1-10$ km/s, with values reaching a few tens of km/s considered significant (Janka & Kresse, 2024). Even models with large neutrino emission asymmetries can not account for natal kicks $\gtrsim 100$ km/s (Burrows et al., 2024b). Based on a single observation of the X-ray-quiescent BH binary VFTS 243, the neutrino natal kick has been estimated to lie around $1-10\ \rm{km/s}$ (Vigna-Gómez et al., 2024), and is consistent with a solution of no natal kick. Thus, if our current understanding of neutrino natal kicks is correct, they have a negligible impact on the binary BH merger time and are highly unlikely to result in a retrograde orbit.
@@
Within the Local Group, several observations have identified high-mass stars and BH binaries at intermediate stages on their path to becoming binary BHs. Notably, the systems Cygnus X-1 (Mirabel & Rodrigues, 2003) and VFTS 243 (Shenar et al., 2022) feature BHs that are candidates for complete collapse. In each case, the BH’s relative orbital velocity exceeds 100 km/s. If these velocities are representative of similar systems, especially just before the formation of the BH, only substantial natal kicks, most likely hydrodynamic in nature, could yield orbital flips. Although there is currently no consensus on the magnitude of BH natal kicks, tentative kinematic evidence suggests that at least some BHs receive kicks above 100 km/s (Nagarajan & El-Badry, 2025).
@@
I briefly outline one scenario in which neutrino natal kicks might influence BHs with stellar companions. Models of isolated binary evolution predict a large population of BHs paired mainly with low-mass companions in the Galaxy (Breivik et al., 2017; Olejak et al., 2020). Many of these systems are expected to originate in wide binaries (Kruckow et al., 2024), making them sensitive to both low and moderate neutrino natal kicks. These wide binaries are likely to be initially eccentric and their spins need not be preferentially aligned, which makes probing orbital flips challenging. Nonetheless, one potential effect is that some wide binaries could be perturbed into tighter, more eccentric orbits, thereby becoming observable by Gaia (Gaia Collaboration et al., 2016). While disentangling the influence of individual wide BH binaries might be difficult, their collective impact could manifest at specific separations—supporting, for example, a super-thermal distribution.
@@ 4 Implication for Spin-Orbit Alignment
Finally, natal kicks have often been considered to affect BH spins—a quantity that is indirectly inferred from gravitational-wave observations of binary BH mergers. In gravitational-wave astronomy, the best-measured spin quantity of binary BHs is given by the effective spin parameter:
@@
where $\chi_{1,2}$ is the magnitude of the dimensionless spin vector, $\cos{\theta_{1,2}}$ is the misalignment angle between the orbital angular momentum vector and the respective spin vectors, and $q=M_{2}/M_{1}\leq 1$ is the mass ratio (e.g., Galaudage et al., 2021, and references therein). The effective spin parameter ranges between $-1\leq\chi_{\rm eff}\leq+1$. It reaches –1 (or +1) for maximally-spinning BHs that are completely anti-aligned (or aligned) with the orbital angular momentum vector. It tends to be close to zero when the spins are small or cancel each other out. For binary BHs formed from isolated binary stars, it has often been assumed that $\chi_{\rm eff}\gtrsim 0$, rarely—if ever—leading to negative values. However, Tauris (2022) showed that randomizing/tossing BH spin axes can explain the observed distribution of the effective spin, alleviating the need for strong kicks. Similarly, Baibhav & Kalogera (2024) finds that relaxing the assumption of spin alignment prior to the second supernova can lead to precessing and retrograde binary BH mergers in isolated environments.
@@
    the spins must be non-zero and aligned,
  2. ii
    there must be no spin tossing during BH formation (i.e. the spin direction must not be random),
@@
    the natal kick needs to be either moderate or large—something that I showed is particularly difficult to reconcile with the complete collapse scenario, and
@@
    the direction of the natal kick need to be either isotropically distributed or biased toward the negative tangential direction.
If the above conditions are met, the binary BH would yield $\chi_{\rm eff}<0$, a value often linked to a dynamical formation scenario, and the time-to-coalescence could be considerably reduced. Our findings suggest that neutrino natal kicks are unlikely to induce orbital flips. If orbital flips do occur, they are more plausibly explained by (larger) hydrodynamic kicks or spin-tossing. Consequently, inferring BH natal kicks from gravitational-wave observations may be challenging. The best prospects could come from electromagnetic observations within the Local Group—specifically, future discoveries of BHs in massive binary systems similar to Cygnus X-1 or VFTS 243. In addition, BHs with luminous companions in wide binaries—potentially detectable by Gaia—may offer valuable insights into BH formation.
@@ Acknowledgements.
I thank Zhanwen Han, Ryosuke Hirai, Hans-Thomas Janka, Jakub Klencki, Mathieu Renzo, Edwin Santiago-Leandro, Taeho Ryu, and Lorenz Zwick for useful discussions and helpful suggestions. I thank Ilya Mandel, Jakob Stegmann, and Thomas Tauris for a careful reading of the manuscript.

### v2

@@ 3 Characteristic orbit of the pre-SN binary
Figure 4 presents $v_{\rm lim}$ for binaries with primary star mass between $1\leq M_{1}/M_{\odot}\leq 1000$ and mass ratios, $q:=M_{2}/M_{1}=\{0.1,1\}$. The black line indicates the $v_{\rm lim}$ threshold that determines which binary BHs (formed without natal kicks) can merge within the age of the Universe; values below this line correspond to binaries that are too wide to merge and therefore cannot become gravitational-wave sources. For a given mass ratio, the region between the blue and red lines corresponds to natal kick magnitudes, which (for that specific orbital configuration) always result in retrograde orbits. Finally, the red line marks the limit above which binaries become gravitationally unbound.
@@
For a star mass of $10\leavevmode\nobreak\ M_{\odot}$ with a mass ratio $q=0.1$ (or $q=1$), the limit orbital speed is $v_{\rm lim}\approx 41$ km/s (or $v_{\rm lim}\approx 213$ km/s). This implies that to disrupt the binary, the natal kicks must exceed roughly $100$ km/s (or $515$ km/s). It is important to note that progenitors of binary BH mergers may exhibit higher orbital speeds, which would in turn raise these limit values. Therefore, our current estimates for the limiting orbital speeds can be considered conservative. In the context of the complete collapse scenario, most models suggest that neutrino natal kicks are $\sim 1-10$ km/s, with values reaching a few tens of km/s considered significant (Janka & Kresse, 2024). Even models with large neutrino emission asymmetries cannot account for natal kicks of $\gtrsim 100$ km/s (Burrows et al., 2024). Based on a single observation of the X-ray-quiescent BH binary VFTS 243, the neutrino natal kick has been estimated to be around $0-10\ \rm{km/s}$ (Vigna-Gómez et al., 2024), consistent with a result of no natal kick. Thus, if our current understanding of neutrino natal kicks is correct, they would have a negligible impact on the binary BH merger time and are highly unlikely to result in a retrograde orbit.
@@
Within the Local Group, several observations have identified high-mass stars and BH binaries at intermediate stages on their path to becoming binary BHs. Notably, the systems Cygnus X-1 (Mirabel & Rodrigues, 2003) and VFTS 243 (Shenar et al., 2022) feature BHs that are candidates for complete collapse. In each case, the relative orbital velocity of the BH exceeds 100 km/s. If these velocities are representative of similar systems, especially just before the formation of the BH, only substantial natal kicks, most likely hydrodynamic in nature, could yield orbital flips. Although there is currently no consensus on the magnitude of BH natal kicks, tentative kinematic evidence suggests that at least some BHs receive kicks above 100 km/s (Nagarajan & El-Badry, 2025).
@@
I briefly outline one scenario where neutrino natal kicks might influence BHs with stellar companions. Models of isolated binary evolution predict a large population of BHs paired mainly with low-mass companions in the Galaxy (Breivik et al., 2017; Olejak et al., 2020). Many of these systems are expected to originate in wide binaries (Kruckow et al., 2024), making them sensitive to both low and moderate neutrino natal kicks. These wide binaries are likely to be initially eccentric and their spins need not be preferentially aligned, which makes probing orbital flips challenging. Nonetheless, one potential effect is that some wide binaries could be perturbed into tighter, more eccentric orbits, thereby becoming observable by Gaia (Gaia Collaboration et al., 2016). While disentangling the influence of individual wide BH binaries might be difficult, their collective impact could be manifested at specific separations; this could support, for example, a super-thermal distribution.
@@ 4 Implication for spin-orbit alignment
Finally, natal kicks have often been considered to affect BH spins: a quantity that is indirectly inferred from gravitational-wave observations of binary BH mergers. In gravitational-wave astronomy, the best-measured spin quantity of binary BHs is given by the effective spin parameter
@@
where $\chi_{1,2}$ is the magnitude of the dimensionless spin vector, $\cos{\theta_{1,2}}$ is the misalignment angle between the orbital angular momentum vector and the respective spin vectors, and $q=M_{2}/M_{1}\leq 1$ is the mass ratio (e.g. Galaudage et al., 2021, and references therein). The effective spin parameter ranges between $1\leq\chi_{\rm eff}\leq+1$. It reaches –1 (or +1) for maximally spinning BHs that are completely anti-aligned (or aligned) with the orbital angular momentum vector. It tends to be close to zero when the spins are small or cancel each other out. For binary BHs formed from isolated binary stars, it has often been assumed that $\chi_{\rm eff}\gtrsim 0$, rarely (if ever) leading to negative values. However, Tauris (2022) showed that randomizing or tossing BH spin axes can explain the observed distribution of the effective spin, alleviating the need for strong kicks. Similarly, Baibhav & Kalogera (2024) finds that relaxing the assumption of spin alignment prior to the second SN can lead to precessing and retrograde binary BH mergers in isolated environments.
@@
    the spins must be non-zero and aligned;
  2. $i$
    there must be no spin tossing during BH formation (i.e. the spin direction must not be random);
@@ 2. $i$
    the natal kick needs to be either moderate or large; this is something that I show is particularly difficult to reconcile with the complete collapse scenario;
@@
    finally, the direction of the natal kick need to be either isotropically distributed or biased toward the negative tangential direction.
If the above conditions are met, the binary BH would yield $\chi_{\rm eff}<0$. This is a value often linked to a dynamical formation scenario and the time to coalescence could be considerably reduced. Our findings suggest that neutrino natal kicks are unlikely to induce orbital flips. If orbital flips do occur, they are more plausibly explained by (larger) hydrodynamic kicks or spin-tossing. Consequently, inferring BH natal kicks from gravitational-wave observations may be challenging. The best prospects could come from electromagnetic observations within the Local Group; specifically, future discoveries of BHs in massive binary systems similar to Cygnus X-1 or VFTS 243. In addition, BHs with luminous companions in wide binaries (potentially detectable by Gaia) could offer valuable insights into BH formation.
@@ Acknowledgements.
I thank Zhanwen Han, Aldana Grichener, Ryosuke Hirai, Hans-Thomas Janka, Jakub Klencki, Mathieu Renzo, Edwin Santiago-Leandro, Taeho Ryu, and Lorenz Zwick for useful discussions and helpful suggestions. I thank Ilya Mandel, Jakob Stegmann, and Thomas M. Tauris for a careful reading of the manuscript. I thank the anonymous referee for their useful suggestions which improved the manuscript.